Dino looks controlled. Parameters are clean again. So it may come back. Lights dropped to 1 and 10. Still over feeding. Scrubber is def working. Cut algae down on it today.
View July 9, 2019 19:22
Three day black out helped control Dino. Added skimmer. Wet vacced sump. Found leak by algae scrubber. Added water with salt to bring salt back up closer to 35. Added 100 micron sock.
View July 4, 2019 02:00
Over fed. Checked basic parameters. Fed anenemes. Still zero phosphate, nitrite, and nitrate.
View June 10, 2019 04:57
Still zero. Added phyto. Double feed. Added salt. Salinity was still low.
View June 6, 2019 14:01
Alex (lionfish) and dom added to sump while Bow is getting fixed (leak)
View June 5, 2019 16:58
Jimmy checked water. Updated to his Hannah info. Found dead skunk shrimp. Head ripped off. Suspect coral banded shrimp.
View June 4, 2019 19:25